### v1

# Universal consistency of the $k$-NN rule in metric spaces and Nagata dimension. III

Abstract.
We prove the last remaining implication allowing to claim the equivalence of the following conditions for a complete separable metric space $X$:
  1. (1)
    The $k$-nearest neighbour classifier is (weakly) universally consistent in $X$,
  2. (2)
    The strong Lebesgue–Besicovitch differentiation property holds in $X$ for every locally finite
    Borel measure,
  3. (3)
    $X$ is sigma-finite dimensional in the sense of Nagata.
[…]
Key words and phrases:
$k$-NN classifier, universal consistency, Nagata dimension, sigma-finite dimensional metric spaces, Lebesgue–Besicovitch property
1991 Mathematics Subject Classification:
62H30, 54F45
1. Introduction
This is the third and last article in a series [4, 11], but it is to a large degree self-contained. The main new result is the following.
{thrm}
Let $\Omega$ be a (complete, separable) metric space that is not sigma-finite dimensional in the sense of Nagata. Then the $k$-NN classifier is not universally consistent in $\Omega$.
Namely, let $(k_{n})$ be an increasing sequence of natural numbers with $k_{n}\in\omega(1)\cap o(n)$. There exist a probability measure $\mu$ on $\Omega$ and a measurable regression function $\eta\colon\Omega\to\{0,1\}$, with $P[\eta(X)=0]=1/2=P[\eta(X)=1]$, as well as a subsequence $(n_{i})$, such that the predictions of the $k$-NN classifier corresponding to the values $n=n_{i}$ and $k=k_{n_{i}}$ converge in probability to the identical value $1$ as $i\to\infty$.
The result holds under any strategy of distance tie-breaking. The proof in essence extends the construction of the measure due to Preiss [17] from the Hilbert space $\ell^{2}$ to any complete separable metric space that is not sigma-finite dimensional in the sense of Nagata, as defined by Preiss elsewhere [18].
There are two consequences of the above result. First, we can now claim the validity of the conjecture put forward in [4].
{thrm}
For a complete separable metric space $\Omega$, the following are equivalent.
  1. (1)
    The $k$-NN classifier is (weakly) universally consistent in $\Omega$.
  2. (2)
    For every locally finite Borel measure $\mu$ on $\Omega$, every $L^{1}(\mu)$-function $f\colon\Omega\to{\mathbb{R}}$ satisies the strong Lebesgue–Besicovitch differentiation property: for $\mu$-a.e. $x\in\Omega$,
    $\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}f(x)\,d\mu(x)\to f(x).$
  3. (3)
    The space $\Omega$ is sigma-finite dimensional in the sense of Nagata.
(Here we assume, like in [3], the uniform random distance tie-breaking of the $k$-NN classifier in case of distance ties.)
[…]
The second consequence is that the weak universal consistency of the $k$-NN rule indeed requires the strong, rather than weak, Lebesgue–Besicovitch differentiation property. Now we can indicate an example of a complete separable metric space and a probability measure on it which satisfies the weak Lebesgue–Besicovitch differentiation property, yet the $k$-NN classifier is not universally consistent. This is the Heisenberg group equipped with a Cygan–Koranyi metric, see e.g. [9, 19, 12] or a relevant discussion and references in [11]. The Heisenberg group is known to have finite metric dimension in the sense of de Groot [5], which, according to Assouad and Quentin de Gromard [1], implies it has the weak Lebesgue–Besicovitch differentiation property for every measure. In [11], we assumed the validity of the result from [3] that the weak Lebesgue–Besicovitch differentiation property suffices for the weak consistency of the $k$-NN classifier, and so we erroneously claimed that the $k$-NN classifier is universally consistent in the Heisenberg group. Our Theorem 1 implies this is not true, as the Heisenberg group is not sigma-finite dimensional in the sense of Nagata [9, 19].
We set things up in Section 2. In Section 3 we discuss the Nagata dimension. The main technical construction of a measure is given in Sect. 4, followed by the proof of Theorem 1 in Sect. 5. Finally, in Section 6 we observe that the weak Lebesgue–Besicovitch property does not in general suffice for the $k$-NN classifier to be consistent.
This note was prompted by a private message [2] from the first author of the paper [3] correcting the statement of their main theorem. For this, I am most grateful to Frédéric Cérou, as well as to László Györfi, Aryeh Kontorovich, Harro Walk, and Roi Weiss for stimulating discussions, without which this paper would not have appeared.
2. Preliminaries
2.1. General setting
We are studying the $k$ nearest neighbour rule for supervised binary classification, working in a complete separable metric space $\Omega$ equipped with a Borel probability measure $\mu$ and a measurable regression function $\eta\colon\Omega\to[0,1]$. (We prefer to denote the “domain” by $\Omega$, reserving the letter $X$ for its random element.) Elements of the random path are sampled in an i.i.d. way following the law $\mu$ and then every point $x$ is labelled 0 or $1$ according to the Bernoulli law with probability of success $\eta(x)$.
[…]
We are interested in the conditons assuring the universal consistency of the $k$-NN classifier, that is, the property that for any pair $\mu,\eta$ as above, as $n,k\to\infty$ so that $k/n\to 0$, the expected misclassification error converges to the Bayes (the smallest possible) error as the size of the sample goes to infinity.
Detailed presentations of the $k$-NN classifier in the metric space setting can be found in numerous sources, including [3, 4, 11].
[…]
2.3. Implication (2) $\Rightarrow$ (1)
This implication follows from the result of Cérou and Guyader [3].
{dfntn}
A complete separable metric space $\Omega$ equipped with a locally finite Borel measure $\mu$ satisfies the strong Lebesgue–Besicovitch differentiation property if for every $L^{1}(\mu)$-function $f\colon\Omega\to{\mathbb{R}}$ and for $\mu$-a.e. $x$,
$\lim_{r\downarrow 0}\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}f(x)\,d\mu(x)=f(x).$ (1)
The metric space $\Omega$ with measure as above satisfies the weak Lebesgue–Besicovitch differentiation property if for every function $f$ as above,
$\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}\eta(x)\,d\mu(x)\to\eta(x)\mbox{ as }r\downarrow 0,$ (2)
where the convergence is in measure, that is, for each ${\epsilon}>0$,
$\mu\left\{x\in\Omega\colon\left|\frac{1}{\mu(B_{r}(x))}\int_{B_{r}(x)}\eta(x)\,d\mu(x)-\eta(x)\right|>{\epsilon}\right\}\to 0\mbox{ when }r\downarrow 0.$
The original theorem in [3] stated that for every complete separable metric space with probability measure $\mu$ satisfying the weak Lebesgue–Besicovitch differentiation property the $k$-NN classifier is consistent with regard to any $\mu$-measurable regression function. However, one of the authors (Frédéric Cérou) has kindly pointed out to us [2] that the very last assertion in the proof of the main result of [3], at the bottom of p. 345, requires the strong Lebesgue–Besicovitch property.
Thus, the corrected form of the result is this.
{thrm}
[Cérou and Guyader [3]] Let $\Omega$ be a complete separable metric space equipped with a probability measure, $\mu$. If $(\Omega,\mu)$ satisfies the strong Lebesgue–Besicovitch differentiation property, then for any $\mu$-measurable regression function $\eta\colon\Omega\to[0,1]$ the $k$-NN classifier is consistent.
[…]
2.4. Implication (3) $\Rightarrow$ (1)
We also note that this implication was established directly in our previous work [4] in the same spirit as the original result of Charles J. Stone for the $k$-NN classifier in the Euclidean space [20].
2.5. Implication (1) $\Rightarrow$ (3)
This implication follows from our Theorem 1.
3. Nagata dimension
This section is largely (though not entirely) a compilation of known results.
3.1. Nagata sigma-finite dimensional spaces
Recall that a family $\gamma$ of subsets of a set $\Omega$ has multiplicity $\leq\delta$ if every intersection of more than $\delta$ different elements of $\gamma$ is empty. In other words,
$\forall x\in\Omega,~~\sum_{V\in\gamma}\chi_{V}(x)\leq\delta,$
where $\chi_{V}$ is the indicator function of $V$.
{dfntn}
A metric subspace $X$ of a metric space $\Omega$ has Nagata dimension $\leq\delta\in{\mathbb{N}}$ on the scale $s>0$ inside of $\Omega$ if every finite family of closed balls in $\Omega$ with centres in $X$ admits a subfamily of multiplicity $\leq\delta+1$ in $\Omega$ which covers all the centres of the original balls. The subspace $X$ has a finite Nagata dimension in $\Omega$ if $X$ has finite dimension in $\Omega$ on some scale $s>0$. Notation for the smallest $\delta$ with the above property: $\dim^{s}_{Nag}(X,\Omega)$ or sometimes simply $\dim_{Nag}(X,\Omega)$.
Following Preiss [18], let us call a family of balls disconnected if the centre of each ball does not belong to any other ball. Here is a mere reformulation of the above definition.
{prpstn}
For a subspace $X$ of a metric space $\Omega$, one has
$\dim^{s}_{Nag}(X,\Omega)\leq\delta$
if and only if every disconnected family of closed balls in $\Omega$ of radii <s<s with centres in $X$ has multiplicity $\leq\delta+1$.
Proof.
Necessity. Let $\gamma$ be a disconnected finite family of closed balls in $\Omega$ of radii <s<s with centres in $X$. Since by assumption $\dim^{s}_{Nag}(X,\Omega)\leq\delta$, the family $\gamma$ admits a subfamily of multiplicity $\leq\delta+1$ covering all the original centres. But the only subfamily that contains all the centres is $\gamma$ itself.
Sufficiency. Let $\gamma$ be a finite family of closed balls in $\Omega$ having radii <s<s and centres in $X$. Denote $C$ the set of centres of those balls. Among all the disconnected subfamilies of $\gamma$ (which exist, e.g., each family containing just one ball is such) choose one, $\gamma^{\prime}$, having the maximal cardinality of the set $C\cap\cup\gamma^{\prime}$. We claim that $C\subseteq\cup\gamma^{\prime}$ (which will finish the argument because the multiplicity of $\gamma^{\prime}$ is bounded by $\delta+1$). Indeed, assume there is a ball, $B\in\gamma$, whose centre, $c\in C$, does not belong to $\cup\gamma^{\prime}$. Remove from $\gamma^{\prime}$ all the balls with centres in $B\cap C$ and add $B$ instead. The new family, $\gamma^{\prime\prime}$, is still disconnected and covers $\left(C\cap\cup\gamma^{\prime}\right)\cup\{c\}$, which contradicts the assumed maximality property of $\gamma^{\prime}$. ∎
In the definition 3.1, as well as in the proposition 3.1, it is convenient to replace closed balls with open ones.
{prpstn}
For a subspace $X$ of a metric space $\Omega$, the following are equivalent.
  1. (1)
    $\dim^{s}_{Nag}(X,\Omega)\leq\delta$,
  2. (2)
    every finite family of open balls in $\Omega$ having radii <s<s with centres in $X$ admits a subfamily of multiplicity $\leq\delta+1$ in $\Omega$ which covers all the centres of the original balls,
  3. (3)
    every disconnected family of open balls in $\Omega$ of radii <s<s with centres in $X$ has multiplicity $\leq\delta+1$.
Proof.
[…]
$\gamma\supseteq\gamma_{1}\supseteq\gamma_{2}\supseteq\ldots\supseteq\gamma_{k}\supseteq\ldots$
with the properties that for each $k$, the family of closed balls $B_{k}$, $B\in\gamma_{k}$ has multiplicity $\leq\delta+1$ in $\Omega$ and covers all the centres of the balls in $\gamma$. Since $\gamma$ is finite, starting with some $k$, the subfamily $\gamma_{k}$ stabilizes, and now it is easy to see that the subfamily $\gamma_{k}$ itself has the multiplicity $\leq\delta+1$, and still covers all the original centres.
[…]
{prpstn}
Let $X$ be a subspace of a metric space $\Omega$, satisfying $\dim^{s}_{Nag}(X,\Omega)\leq\delta$. Then $\dim^{s}_{Nag}(\bar{X},\Omega)\leq\delta$, where $\bar{X}$ is the closure of $X$ in $\Omega$.
Proof.
Let $\gamma$ be a finite disconnected family of open balls in $\Omega$ of radii <s<s, centred in $\bar{X}$. Let $y\in\Omega$, and let $\gamma^{\prime}$ consist of all balls in $\gamma$ containing $y$. Choose ${\varepsilon}>0$ so small that the open $\varepsilon$}-ball around $y$ is contained in every element of $\gamma^{\prime}$. For every open ball $B\in\gamma^{\prime}$, denote $y_{B}$ the centre and $r_{B}$ the radius. We can additionally assume that ${\varepsilon}<r_{B}$. Denote $B^{\prime}$ an open ball of radius $r_{B}-{\varepsilon}>0$, centred at a point $x_{B}\in X$ satisfying $d(x_{B},y_{B})<{\varepsilon}/2$. Note that $y\in B^{\prime}$ by the triangle inequality. Again by the triangle inequality, the family $\{B^{\prime}\colon B\in\gamma^{\prime}\}$ is disconnected, and it has radii <s<s. Therefore, $y$ only belongs to $\leq\delta+1$ balls $B^{\prime}$, $B\in\gamma^{\prime}$. Consequently, the cardinality of $\gamma^{\prime}$ is bounded by $\delta+1$. ∎
{prpstn}
If $X$ and $Y$ are two subspaces of a metric space $\Omega$, having finite Nagata dimension in $\Omega$ on the scales $s_{1}$ and $s_{2}$ respectively, then $X\cup Y$ has a finite Nagata dimension in $\Omega$, with $\dim_{Nag}(X\cup Y,\Omega)\leq\dim_{Nag}(X,\Omega)+\dim_{Nag}(Y,\Omega)$, on the scale $\min\{s_{1},s_{2}\}$.
Proof.
Given a finite family of balls $\gamma$ in $\Omega$ of radii $\min\{s_{1},s_{2}\}$ centred in elements of $X\cup Y$, represent it as $\gamma=\gamma_{X}\cup\gamma_{Y}$, where the balls in $\gamma_{X}$ are centered in $X$, and the balls in $\gamma_{Y}$ are centred in $Y$. Select a subfamily $\gamma_{X}^{\prime}$ of multiplicity $\dim_{Nag}(X,\Omega)$ containing all centres of balls in $\gamma_{X}$, and a subfamily $\gamma_{Y}^{\prime}$ of multiplicity $\dim_{Nag}(Y,\Omega)$ containing all centres of balls in $\gamma_{Y}$, and form their union. ∎
{dfntn}
[Preiss [18]] A metric space $\Omega$ is sigma-finite dimensional in the sense of Nagata if $\Omega=\cup_{i=1}^{\infty}X_{n}$, where every subspace $X_{n}$ has finite Nagata dimension in $\Omega$ on some scale $s_{n}>0$ (where the scales $s_{n}$ are possibly all different).
{rmrk}
Due to Proposition 3.1, in the above definition we can assume the subspaces $X_{n}$ to be closed in $\Omega$. By Proposition 3.1, we can assume $X_{n}$ to form an increasing chain.
{prpstn}
Every separable metric space $\Omega$ that is not sigma-finite dimensional (in itself) in the sense of Nagata contains a non-empty closed subspace $\Upsilon$ none of whose non-empty relatively open subsets has finite Nagata dimension in $\Omega$.
Proof.
Let us say, slightly extending the definition 3.1 of Preiss, that a subset $A$ of $\Omega$ is sigma-finite dimensional in $\Omega$ if one can represent $A$ as the union of a countable family of subsets $A_{n}$ each of which is finite-dimensional in $\Omega$ on some positive scale.
Denote $\gamma$ the family of all non-empty open subsets of $\Omega$ that are sigma-finite dimensional in $\Omega$. Let $O$ be the union of $\gamma$. If it happens that $O=\emptyset$, there is nothing left to prove, so we may as well assume $\gamma\neq\emptyset$. Since $O$ is a separable metric space, it is Lindelöf, and one can select a countable subfamily of $\gamma$ whose union is $O$ ([8], Thm. 3.8.1). It follows that $O$ itself is sigma-finite dimensional in $\Omega$, so it is the largest open subset sigma-finite dimensional in $\Omega$.
Denote $\Upsilon=\Omega\setminus O$. Since $\Omega$ is not sigma-finite dimensional in itself, the subspace $\Upsilon$ is non-empty, and it is also closed. Let now $V$ be a non-empty relatively open subset of $\Upsilon$. Thus, there exists an open $U\subseteq\Omega$ with $U\cap\Upsilon=V$. Suppose that $V$ is sigma-finite dimensional in $X$. Since $Y$ is closed, $U\setminus V=U\setminus Y$ is an open subset of $\Omega$, in fact of $O$, hence $U\setminus V$ is sigma-finite dimensional in $\Omega$. It follows that $U=V\cup(U\setminus V)$ is an open sigma-finite dimensional subset of $\Omega$ that is not contained in $O$, in contradiction to the maximality of $O$ with these properties. We conclude: every non-empty relatively open subset of $\Upsilon$ is not sigma-finite dimensional in $\Omega$, in particular, has infinite Nagata dimension in $\Omega$. ∎
{rmrk}
In [18], Preiss seems to imply (p. 567, lines 3-4) that every complete separable metric space $\Omega$ that is not sigma-finite dimensional (in itself) in the sense of Nagata contains a non-empty closed subspace $\Upsilon$ containing no non-empty (relatively) open subsets of finite Nagata dimension in $\Upsilon$ proper, but I am unable to verify this claim. The purported proof of it that my student Sushma Kumari and I had come up with and included in her Ph.D. thesis (Lemma 1.4.2, p. 35 in [10]) appears deficient.
3.2. Further remarks
{rmrk}
A metric space is said to have finite Nagata dimension if it has finite Nagata dimension in itself. The Nagata–Ostrand theorem [13, 14, 16] states that the Lebesgue covering dimension of a metrizable topological space is the smallest Nagata dimension of a compatible metric on the space (and in fact this is true on every scale $s>0$). This is the historical origin of the concept of the Nagata metric dimension.
{xmpl}
A metric space has Nagata dimension zero on the scale $\infty$ if and only if it is non-archimedian, that is, satisfies the strong triangle inequality, $d(x,z)\leq\max\{d(x,y),d(y,z)\}$.
{xmpl}
It is easy to see, especially using Prop. 3.1, that $\dim_{Nag}({\mathbb{R}})=1$.
In the context of statistical learning theory, Nagata dimension captures in an abstract setting the geometry behind Stone’s lemma used in the proof of the universal consistency of the $k$-NN classifier in the finite-dimensional Euclidean space (see Sect. 5.3 in [7]).
{xmpl}
The Nagata dimension of the Euclidean space $\ell^{2}(d)$ is finite, and it is bounded by $\gamma_{d}-1$, where $\gamma_{d}$ is the value of the constant in Stone’s geometric lemma (ibid.).
Indeed, let $x_{1},\ldots,x_{\gamma_{d}+1}$ be points belonging to a ball with centre $x$. Using the argument in the proof of Stone’s geometric lemma with $k=1$, mark $\leq\gamma_{d}$ points $x_{i}$ belonging to the $\leq\gamma_{d}$ cones with apex at $x$. At least one point, say $x_{j}$, has not been marked; it belongs to some cone, which therefore already contains a marked point, say $x_{i}$, different from $x_{j}$, and ${\left\|\,x_{i}-x_{j}\,\right\|}\leq{\left\|\,x_{j}-x\,\right\|}$.
{rmrk}
In ${\mathbb{R}}^{2}={\mathbb{C}}$ the family of closed balls of radius one centred at the vectors $\exp(2\pi ki/5)$, $k=1,2,\ldots,5$, has multiplicity $5$ and admits no proper subfamily containing all the centres. Therefore, the Nagata dimension of $\ell^{2}(2)$ is at least $5$. Since the plane can be covered with $6$ cones having the central angle $\pi/3$, Example 3.2 implies that $\dim_{Nag}(\ell^{2}(2))=5$.
{rmrk}
The exact values of the Nagata dimension of the higher-dimensional Euclidean spaces seem to be still unknown, this was stated as an open problem in [15].
{rmrk}
A very good reference to a great variety of distance dimensions, including the Nagata dimension, even in the absence of triangle inequality, and their applications to measure differentiation, is the article [1].
4. The measure construction
4.1. The indexing tree
Let $\Omega$ be a separable complete metric space that is not sigma-finite dimensional in the sense of Nagata. Fix a closed subspace $\Upsilon\subseteq\Omega$ with the property that none of its non-empty relatively open subsets has finite Nagata dimension in $\Omega$ (Proposition 3.1).
We will work with an unbounded sequence $(m_{i})_{i=0}^{\infty}$ of positive natural numbers starting with $m_{0}=1$, which is arbitrary at this stage, but will be recursively specified at a later point. The only condition we impose from the start is that $m_{i}\geq 2$ for $i\geq 1$. Form a rooted prefix tree, $T$, with vertices being finite words $t=t_{1}t_{2}\ldots t_{i}$, $i\in{\mathbb{N}}$, $t_{j}\in[m_{j}]=\{1,2,\ldots,m_{j}\}$, $j=1,2,\ldots,i$. For such a $t$, we denote $i=\lvert t\rvert$. The root is the empty word, denoted $\ast$; so, $\lvert\ast\rvert=0$. In this way, the children of the node $t$ with $\lvert t\rvert=i$ are the words $t1,t2,\ldots,tm_{i+1}$.
4.2. Selecting points $x^{t},y^{t}$
We will now select two families of points of the metric space $\Omega$ indexed with the vertices of the tree $T$, namely, $x^{t}\in\Omega$ for all $t\in T$, and $y^{t}\in\Upsilon$ for $t\in T\setminus\{\ast\}$, as well as numbers $0<{\epsilon}^{t}<2^{-\lvert t\rvert}$, $t\in T$, with the following properties for each $t\in T$, $\lvert t\rvert=i$, and every $j=1,2,\ldots,m_{i+1}$, $\ell=1,2,\ldots,m_{i+2}$:
  1. (1)
    for every $z\in\bar{B}_{{\epsilon}^{t}}(y^{tj})$, the point $x^{t}$ is strictly closer to $z$ than any point contained in any of the balls $\bar{B}_{{\epsilon}^{t}}(y^{ts})$, $s\neq j$,
  2. (2)
    $x^{tj}\in B_{{\epsilon}^{t}}(y^{tj})$, and
  3. (3)
    $B_{{\epsilon}^{tj}}(y^{tj\ell})\subseteq B_{{\epsilon}^{t}}(y^{tj})$.
The existence of such points is a direct consequence of our assumptions on $\Omega$ and $\Upsilon$, and we will do it in some detail.
4.2.1. Base of recursion
At the first step, we use the fact that $\Upsilon$ itself is not Nagata finite dimensional in $\Omega$. In particular, as $\dim^{s}_{Nag}(\Upsilon,\Omega)>m_{1}$ on any scale $s>0$, there exists a disconnected collection of $m_{1}$ balls of radii <s<s centred in $\Upsilon$ (which balls we can take open by Proposition 3.1), having a point in $\Omega$ in common. At the base step, the choice of a scale does not matter, we can take $s=1$.
Denote the centres of the selected balls $y^{1},y^{2},\ldots,y^{m_{1}}\in\Upsilon$, the radii $r_{1}^{1},\ldots,r_{1}^{m_{1}}>0$, and the common point of the open balls $x^{\ast}\in\Omega$. Take ${\epsilon}^{\ast}>0$ so small that ${\epsilon}^{\ast}<1$ and for each $1,2,\ldots,m_{1}, iji\neq j$, we have
${\epsilon}^{\ast}<\frac{1}{4}\left(d(y^{i},y^{j})-d(y^{i},x^{\ast})\right).$
It follows now by the triangle inequality that for any point, $z$, in the closed ball $\bar{B}_{{\epsilon}^{\ast}}(y^{i})$ the point $x^{\ast}$ is strictly closer to $z$ than is any point, $w$, of any closed ball $\bar{B}_{{\epsilon}^{\ast}}(y^{j})$, $j\neq i$:
$d(z,w)-d(z,x^{\ast})$ $\geq d(y^{i},y^{j})-d(y^{i},x^{\ast})-4{\epsilon}^{\ast}$
$>0.$
4.2.2. Step of recursion
Now, let $i\geq 1$ and suppose that $x^{t}\in\Omega$ and ${\epsilon}^{t}>0$ have been selected for all $t$ with $0\leq\lvert t\rvert<i$, and $y^{t}\in\Upsilon$ have been chosen for all $t$ with $1\leq\lvert t\rvert\leq i$, in a way to satisfy the properties (1)-(3) in subs. 4.2.
Fix a $t$ with $\lvert t\rvert=i$ and denote $t^{-}$ the (i1)(i-1)-prefix of $t$, that is, $t$ with the last letter removed. We now have to select $x^{t}\in\Omega$ and $y^{tj}\in\Upsilon$, $j=1,2,\ldots,m_{i+1}$, all within the open ball $B_{{\epsilon}^{t^{-}}}(y^{t})$, as well as the value $0<{\epsilon}^{t}<2^{-i}$.
The relative open ball $B_{{\epsilon}^{t^{-}}/2}(y^{t})\cap\Upsilon$ is non-empty (it contains $y^{t}$) and so is not finite dimensional in $\Omega$ in the sense of Nagata by our choice of $\Upsilon$, in particular it is infinite dimensional in $\Omega$ on all scales, including $s={\epsilon}^{t^{-}}/2$. There exist points $y^{t1},y^{t2},\ldots,y^{tm_{i+1}}\in B_{{\epsilon}^{t^{-}}/2}(y^{t})\cap\Upsilon$ and radii $0<r^{t1},r^{t2},\ldots,r^{tm_{i+1}}<{\epsilon}^{t^{-}}/2$ such that the open balls $B_{r^{t1}}(y^{t1}),B_{r^{t2}}(y^{t2}),\ldots,B_{r^{tm_{i+1}}}(y^{tm_{i+1}})$ are disconnected (do not contain each other’s centres) and have a point in $\Omega$ in common, which point we denote $x^{t}$. By the triangle inequality, all open balls $B_{r^{tj}}(y^{tj})$, $j\in[m_{i+1}]$ are contained in $B_{{\epsilon}^{t^{-}}}(y^{t})$. In particular, $x^{t}\in B_{{\epsilon}^{t^{-}}}(y^{t})$. Just like at the first step of recursion, we can find ${\epsilon}^{t}\in(0,2^{-i})$ with the property that for every $j\in[m_{i+1}]$ and every point $z\in\bar{B}_{{\epsilon}^{t}}(y^{tj})$, the point $x^{t}$ is strictly closer to $z$ that any point of any closed ball $\bar{B}_{{\epsilon}^{t}}(y^{ts})$, $s\neq j$. The recursion step is now completed.
4.3. Compact set $K$ supporting two measures
By the condition 4.2(3), for any $t$ with $\lvert t\rvert=i\geq 1$ and any word $s$, we get $d(y^{t},y^{ts})<2^{-i+1}$. Also, the condition 4.2(2) says that $d(y^{t},x^{t})<2^{-i+1}$. Consequently, all the points $y^{t}$, $x^{t}$ with $\lvert t\rvert\geq i$ are contained within the distance $2^{-i+2}$ from the finite set of points $y^{t}$, $\lvert t\rvert=i$. This implies the entire countable set of points $x^{t}$, $\lvert t\rvert\geq 0$, $y^{t}$, $\lvert t\rvert\geq 1$ is precompact. Denote $K$ its closure, a compact subset of $\Omega$.
We will now construct two finite measures supported on $K$, each one of total mass $1/2$: a purely atomic measure $\mu_{1}$, having $x^{t}$, $t\in T$ as atoms, and a non-atomic measure, $\mu_{0}$, which is a weak limit of the sequence of uniform measures supported on the points $y^{t}$, $\lvert t\rvert=i$, as $i\to\infty$.
4.3.1. Measure $\mu_{1}$
Fix a sequence $(\gamma_{n})_{n=0}^{\infty}$, $\gamma_{n}>0$, with $\sum_{n=0}^{\infty}\gamma_{n}=1/2$. Now define for each $t\in T$
$\mu_{1}\{x^{t}\}=\frac{\gamma_{\lvert t\rvert}}{\prod_{0\leq j\leq\lvert t\rvert}m_{j}}.$ (3)
Notice that for each $i=0,1,\ldots$,
$\left|\{x^{t}\colon\lvert t\rvert=i\}\right|=\prod_{0\leq j\leq i}m_{j},$
and so $\mu_{1}\{x^{t}\colon\lvert t\rvert=i\}=\gamma_{i}$, and $\sum_{t\in T}\mu_{1}\{x^{t}\}=\mu_{1}(K)=1/2$.
4.3.2. Measure $\mu_{0}$
For every $i=1,2,\ldots$ let $\mu_{0}^{(i)}$ be a uniform measure of total mass $1/2$ supported on the points $y^{t}$ with $\lvert t\rvert=i$, that is,
$\mu^{(i)}\{y^{t}\}=\frac{1}{2\prod_{1\leq j\leq i}m_{j}}.$
Since for each $i=1,2,\ldots$,
$\left|\{y^{t}\colon\lvert t\rvert=i\}\right|=\prod_{1\leq j\leq i}m_{j},$
we have $\mu^{(i)}_{0}(K)=1/2$ for each $i$.
The sequence of measures $(\mu_{0}^{(i)})$ converges in the weak topology in the space of finite measures on the compact space $K$. Indeed, the weak topology (that is, the weak topology on the dual space of $C(K)$) on the measures of fixed mass on a compact space is metrized by the mass transportation distance (see e.g. [21], corollary 6.13). For every $t$, $\lvert t\rvert\geq 1$, the $\mu_{0}^{(i)}$-measure of the ball $\bar{B}_{{\epsilon}^{t^{-}}}(y^{t})$ is equal to $1/(2\prod_{j\leq i}m_{i})$, because the ball only contains exactly one point among all $y^{t}$ with $\lvert t\rvert=i$, namely, its centre. If now $j>i$, then the number of points $y^{s}$ with $\lvert s\rvert=j$ contained in the same ball $\bar{B}_{{\epsilon}^{t^{-}}}(y^{t})$ is equal to $\prod_{\ell=i+1}^{j}m_{\ell}$, and so the total mass of the ball under the measure $\mu_{0}^{(j)}$ is the same as under $\mu_{0}^{(i)}$. The transport of mass between the restrictions of the two measures on the ball requires shifting one of them by a distance $\leq{\epsilon}^{t^{-}}<2^{-i}$. We conclude that the overall mass transportation distance between $\mu_{0}^{(i)}$ and $\mu_{0}^{(j)}$ is bounded by $2^{-i}$ times the total mass, $1/2$. This way, the sequence $(\mu_{0}^{(i)})$ is Cauchy, hence convergent in the weak topology on the compact space of measures of mass $1/2$. Denote the limit $\mu_{0}$. It also satisfies $\mu_{0}(K)=1/2$.
For every $i$, the union of the closed balls $\bar{B}_{{\epsilon}^{t^{-}}}(y^{t})$, $\lvert t\rvert=i$ has full $\mu_{0}$-measure, $1/2$. In particular, it follows that the measure $\mu_{0}$ is diffuse, because for each $t$ with $\lvert t\rvert=i$, it has the property $\mu_{0}(\bar{B}_{{\epsilon}^{t^{-}}}(y^{t}))=1/(2\prod_{0\leq j\leq i}m_{i})\to 0$ as $i\to\infty$.
4.3.3. The probability measure $\mu$ and the regression function $\eta$
Since $\mu_{1}$ is purely atomic and $\mu_{0}$ is diffuse, the two measures are mutually singular. Define a probability measure $\mu=\mu_{0}+\mu_{1}$. Since $\mu_{0}\perp\mu_{1}$, there is a measurable function $\eta\colon X\to\{0,1\}$ that is $\mu_{0}$-a.e. equal to 0 e $\mu_{1}$-a.e. equal to $1$. (The indicator function of the set $\{x^{t}\colon t\in T\}$ is such.)
This $(\mu,\eta)$ will become our learning problem. The Bayes (optimal) classifier for this problem is a.e. equal to the binary function $\eta$, so will assume the value $1$ with probability half. As we will see, under a suitable selection of the sequence $(m_{i})$ used to construct $\mu$ and $\eta$, the $k$-NN classifier with a given sequence of values $(k_{n})_{n=1}^{\infty}$ does not asymptotically converge to $\eta$ and so is not consistent under this problem.
5. Inconsistency of the $k$-NN under $(\mu,\eta)$
5.1. Selecting $(m_{i})$ and $(n_{i})$
Let the sequence $(k_{n})_{n=1}^{\infty}$ of values of $k$ be given, satisfying $k_{n}\to\infty$ and $k_{n}/n\to 0$. Fix a summable sequence of values of the risk $\delta_{i}>0,i\in{\mathbb{N}}$. Remember that $m_{0}=1$. This is the base of the recursion.
5.1.1. Step $m_{i}\to n_{i}$
Let $i=0,1,\ldots$, and suppose that the values $m_{j}$, $j=1,2,\ldots,i$ have already been selected. We want to choose $n_{i}$ so large that, with confidence $1-\delta_{i}$, each of the points $x^{t}$, $\lvert t\rvert=i$ appears among the i.i.d. random elements $X_{1},\ldots,X_{n_{i}}$, following the law $\mu$ as above, no fewer than $k_{n_{i}}$ times. Remembering that the mass of $x^{t}$ is given by eq. (3) and that $k_{n}/n\to 0$, first assume that $n$ is so large that
$\frac{k_{n}}{n}<\frac{1}{2}\mu\{x^{t}\}=\frac{1}{2}\mu_{1}\{x^{t}\}=\frac{\gamma_{i}}{2\prod_{0\leq j\leq i}m_{j}}.$ (4)
By the Chernoff bound, the probability of the empirical measure of the singleton $\{x^{t}\}$ deviating from its expectation, $\mu\{x^{t}\}$, by more than half its value is limited by
$\exp\left[{-2n\left(\frac{\gamma_{i}}{2\prod_{0\leq j\leq i}m_{i}}\right)^{2}}\right].$
The probability that such a deviation happens for at least one among the points $x^{t}$, $\lvert t\rvert=i$, is limited by the union bound by
$\prod_{0\leq j\leq i}m_{j}\exp\left[{-2n\left(\frac{\gamma_{i}}{2\prod_{0\leq j\leq i}m_{j}}\right)^{2}}\right].$
We want the above to be limited by $\delta_{i}$. Taking the natural logarithm of the inequality, we arrive at
$\frac{\gamma^{2}_{i}n}{2\prod_{0\leq j\leq i}m^{2}_{j}}>\sum_{0\leq j\leq i}\log m_{j}-\log\delta_{i},$
that is,
$n>\frac{2\prod_{0\leq j\leq i}m^{2}_{j}}{\gamma^{2}_{i}}\left(\sum_{0\leq j\leq i}\log m_{j}-\log\delta_{i}\right).$
Choose any value $n=n_{i}$ satisfying the last inequality and such that eq. (4) holds for $n\geq n_{i}$. If now $n\geq n_{i}$, then, with confidence $1-\delta_{i}$, the empirical measure of every point $x^{t}$, $\lvert t\rvert=i$ will be at least $(1/2)\mu\{x^{t}\}>k_{n_{i}}/n_{i}$. Consequently, each point $x^{t}$, $\lvert t\rvert=i$ will appear at least $k_{n}$ times among the members of a random i.i.d. $n$-sample following the law $\mu$.
5.1.2. Step $n_{i}\to m_{i+1}$
Now we will choose $m_{i+1}$. We want this number to be so large that the proportion of the number of balls $B_{{\epsilon}^{t}}(y^{tj})$, $\lvert t\rvert=i$, $j=1,2,\ldots,m_{i+1}$, containing at least $k_{n_{i}}/2$ elements of a random i.i.d. $n_{i}$-sample following the law $\mu$ is less than $\delta_{i}$. The maximal possible number of such balls is limited by $2n_{i}/k_{n_{i}}$. So it is enough to require that the ratio of $2n_{i}/k_{n_{i}}$ to the total number of points $y^{t}$ with $\lvert t\rvert=i+1$ (that is, the number of balls under consideration) is less than $\delta_{i}$:
$\frac{2n_{i}}{k_{n_{i}}}<\delta_{i}\prod_{j=0}^{i+1}m_{j}=\delta_{i}\left(\prod_{j=0}^{i}m_{j}\right)m_{i+1}.$
This means selecting $m_{i+1}$ with
$m_{i+1}>\frac{2n_{i}}{k_{n_{i}}\delta_{i}\prod_{j=0}^{i}m_{j}}.$
Now the recursive step is accomplished.
Notice in passing that as $n_{i}/\prod_{j=0}^{i}m_{j}\geq k_{n_{i}}$ by eq. (4), the r.h.s. of the above expression is no less than $2/\delta_{i}$, and so $m_{i}\to\infty$. (This assures that the measure $\mu_{0}$ is diffuse, see the end of Subs. 4.3.2, though for that it would suffice that each $m_{i}\geq 2$.)
5.2. Absence of weak consistency of the $k$-NN rule under $(\mu,\eta)$
5.2.1. Predictor for $x^{t}$
Each point $x^{t}$ is an atom, labelled deterministically $1$. Under such conditions, it is well known and easily proved that the label predicted by the $k$-NN classifier in the limit $n\to\infty$ converges on the set of atoms in probability to $1$.
It remains to analyse the situation with $x\in K$ that is not of the form $x^{t}$ and thus will $\mu$-almost surely be labelled 0. We will show that the label predicted by the $k$-NN classifier does not converge in probability to zero.
5.2.2. $x\in K\setminus\{x^{t}\colon t\in T\}$
Fix $i\in{\mathbb{N}}$, $i\geq 1$ throughout the subsection.
Going back to Subs. 4.3, we notice that, since each point $y^{t}$ has zero measure, $\mu$-almost all points $x$ of the set $K\setminus\{x^{t}\colon t\in T\}$ (in fact, all except a finite number) are contained in the set $\cup_{\lvert t\rvert=i}\cup_{j=1}^{m_{i+1}}B_{{\epsilon}^{t}}(y^{tj})$. Since the balls are pairwise disjoint, for $\mu$-almost all $x$ there are unique $t\in T$ with $\lvert t\rvert=i$ and $j=1,2,\ldots,m_{i+1}$ such that $x\in B_{{\epsilon}^{t}}(y^{tj})$.
Now, $x^{t}$ is strictly closer to $x$ than any point of any other ball of the same form, $B_{{\epsilon}^{t}}(y^{ts})$, $s\neq j$ (property 4.2(1)). If we form a closed ball around $x$ of radius $d(x,x^{t})$, the intersection of this ball with $K$ will be the same as the intersection of the ball $B_{{\epsilon}^{t}}(y^{tj})$ with $K$ plus $\{x^{t}\}$.
Hence, with probability $1-\delta_{i}$, conditionally on $X\notin\{x^{t}\colon t\in T\}$, the open ball around a random element $X\sim\mu$ having the same radius as the distance to the nearest $x^{t}$, $\lvert t\rvert=i$, contains strictly fewer than $k_{n_{i}}/2$ elements of the random i.i.d. $n_{i}$-sample following the law $\mu$ (Subs. 5.1.2).
At the same time, with probability $1-\delta_{i}$, each of the points $x^{t}$, $\lvert t\rvert=i$ appears at least $k_{n_{i}}$ times among the elements of the random i.i.d. $n_{i}$-sample following the law $\mu$ (Subs. 5.1).
We conclude: with probability $1-2\delta_{i}$, strictly more than half of the $k=k_{n_{i}}$ nearest neighbours to $X$ among the random i.i.d. $n_{i}$-sample are equal to one of the points $x^{t}$, $\lvert t\rvert=i$, namely the one nearest to $X$. And the points $x^{t}$ are almost surely labelled $1$.
This means that, with probability $1-2\delta_{i}$, a random element $X\notin\{x^{t}\colon t\in T\}$ will get labelled $1$ by the $k$-NN classifier at the step $(n_{i},k_{n_{i}})$.
5.2.3. As $i\to\infty$
As $\sum\delta_{i}<\infty$, by Borel–Cantelli, the subsequence of $k$-NN classifiers corresponding to $(n_{i},k_{n_{i}})$ converges to the identical value $1$ over $\Omega\setminus\{x^{t}\colon t\in T\}$ almost surely, in particular, in probability. Combining this with the observation in Subs. 5.2.1, we conclude: the subsequence of $k$-NN classifiers corresponding to $(n_{i},k_{n_{i}})$ converges to the identical value $1$ over all of $\Omega$ in probability.
This implies that the sequence of $k$-NN classifiers corresponding to the original sequence $(n,k_{n})$ does not converge in probability to the Bayes classifier, $\eta$, and so the $k$-NN rule is not weakly consistent under $(\mu,\eta)$.
6. Weak Lebesgue–Besicovitch property is not sufficient for the consistency of the $k$-NN classifier
Here we aim to correct a statement made in the previous article [11], namely, Corollary 3.18 in Section 3.
In that section, we reminded the weaker notion of dimension of a metric space in the sense of de Groot [5], which is always less than or equal to the Nagata dimension. According to a result of Assouad and Quentin de Gromard ([1], Prop. 3.3.1(b)+Prop. 3.1), a complete separable metric space having finite de Groot dimension satisies the weak Lebesgue–Besicovitch differentiation property for every locally finite Borel measure. A classical example of a metric space of finite de Groot dimension (moreover, having a doubling property) is the Heisenberg group $\mathbb{H}$}, equipped with the Cygan-Korányi metric [6]. At the same time, as shown, independently, by Korányi and Reimann ([9], p. 17) and Sawyer and Wheeden ([19], Lemma 4.4, p. 863), the Nagata dimension of the Heisenberg group $\mathbb{H}$} is infinite, and in fact their argument shows that $\mathbb{H}$} is not sigma-finite dimensional in the sense of Nagata ([11], Corol. 3.20). For a detailed presentation of all of this, with proofs and references, we refer the reader to Sect. 3 in [11].
Now, in [11], we combined the above with the main result as stated in [3] to claim that the $k$-NN classifier in the Heisenberg group is universally consistent. The main result of the present paper, Thm. 1, shows this not to be true.
Since the Heisenberg group $\mathbb{H}$}, equipped with the Cygan-Korányi metric, is not sigma-finite dimensional in the sense of Nagata by [9, 19], the $k$-NN classifier is not weakly consistent in it – even if, according to [1], the group $\mathbb{H}$} with the Cygan-Korányi metric satisfies the weak Lebesgue–Besicovitch differentiation property for every locally finite measure. Modulo our measure construction in Section 4, this gives an example of a complete separable metric space equipped with a probability measure that satisfies the weak Lebesgue–Besicovitch differentiation property but in which the $k$-NN classifier is not consistent.